# The algebraic structure of hyperbolic graph braid groups

### 13 Mar 2024 v1

@@ Abstract.
Genevois has recently classified which graph braid groups on $\geq 3$ strands are word hyperbolic. In the $3$-strand case, he asked whether all such word hyperbolic group are actually free; this reduced to checking two infinite classes of graphs. We give positive and negative answers to this question: one class of graphs always has free $3$-strand braid group, while the braid groups in the other class usually contain surface subgroups.
@@ 1.2. Hyperbolic graph braid groups
On the other hand, it turns out that the $3$-strand graph braid group of a pulsar graph is almost never free. In fact it was already known that $\operatorname{Conf}_{3}(\Theta_{4})$ is homotopy equivalent to a closed, orientable surface of genus $3$ [CL18, WG19]. We provide a simple proof of this fact. Specifically, we show explicitly that with the simplest admissible subdivision of $\Theta_{4}$, the space $\mathrm{Conf}^{\square}_{3}({\Theta_{4}})$ is homeomorphic to a closed, orientable surface of genus $3$. A consequence of this is that $\mathrm{Conf}^{\square}_{3}({\mathcal{P}})$ has non-trivial second homology for any pulsar graph $\mathcal{P}$ containing $\Theta_{4}$. This proves:
@@
If $\mathcal{P}_{m,n_{1},n_{2}}$ is a pulsar graph with $m\geq 4$ (in the notation of Definition 4.1), then $B_{3}(\mathcal{P}_{m,n_{1},n_{2}})$ is not free.
@@
In particular, $\chi(\operatorname{Conf}_{3}(\Theta_{7}))=0$ and $\chi(\operatorname{Conf}_{3}(\Theta_{m}))>0$ when $m>7$. For $m>7$, this implies that $B_{3}(\Theta_{m})$ is not the fundamental group of an aspherical $3$-manifold (a doubling argument shows these have nonpositive Euler characteristic). This answers a part of Question 1.2
Finally, we obtain the following free product decomposition for the 3-strand graph braid groups of pulsar graphs:
Theorem 1.6.
This again has implications for Question 1.2, as it shows that $B_{3}(\mathcal{P})$ is not one-ended when $\mathcal{P}$ has a non-trivial ray, i.e. when it is not a generalized theta graph.
Our main tool to prove Theorems 1.3 and 1.6 is the PL Morse theory of Bestvina–Brady  [BB97]. This was inspired by Bestvina’s proof that braid groups of the tripod graph are free [Bes08, Section 4]. Discrete notions of Morse theory have been used to study graph braid groups before, however most of the earlier papers use Forman’s version [For98].
@@ 4.3. Free product decomposition
In this section we show that if $\mathcal{P}_{m,n_{1},n_{2}}$ is the pulsar graph from Definition 4.1, then the group $B_{3}(\mathcal{P}_{m,n_{1},n_{2}})$ decomposes as a free product of the form $B_{3}(\Theta_{m})*\mathbb{F}$, where $\mathbb{F}$ is a free group. This proves Thoerem 1.6 of the introduction.

### 21 Mar 2024 v2

@@ Abstract.
Genevois recently classified which graph braid groups on $\geq 3$ strands are word hyperbolic. In the $3$-strand case, he asked whether all such word hyperbolic groups are actually free; this reduced to checking two infinite classes of graphs: sun and pulsar graphs. We prove that $3$-strand braid groups of sun graphs are free. On the other hand, it was known to experts that $3$-strand braid groups of most pulsar graphs contain surface subgroups. We provide a simple proof of this and prove an additional structure theorem for these groups.
@@ 1.2. Hyperbolic graph braid groups
On the other hand, it turns out that the $3$-strand graph braid group of a pulsar graph is almost never free. In fact it was already known that $\operatorname{Conf}_{3}(\Theta_{4})$ is homotopy equivalent to a closed, orientable surface of genus $3$ [CL18, WG19]. We provide a simple proof of this: we show that with the simplest admissible subdivision of $\Theta_{4}$, the space $\mathrm{Conf}^{\square}_{3}({\Theta_{4}})$ is homeomorphic to this surface. Since $B_{3}(\Theta_{4})$ injects into $B_{3}(\mathcal{P}_{m,n_{1},n_{2}})$ by [Abr00, Corollary 3.14], it follows that $B_{3}(\mathcal{P}_{m,n_{1},n_{2}})$ is not free for $m\geq 4$.
In this case, we still prove some additional structure for the 3-strand graph braid groups of pulsar graphs:
@@
This still has implications for Question 1.2, as it shows that $B_{3}(\mathcal{P})$ is not one-ended when $\mathcal{P}$ has a non-trivial ray, i.e. when it is not a generalized theta graph.
@@
In particular, $\chi(\operatorname{Conf}_{3}(\Theta_{7}))=0$ and $\chi(\operatorname{Conf}_{3}(\Theta_{m}))>0$ when $m>7$. For $m>7$, this implies that $B_{3}(\Theta_{m})$ is not the fundamental group of an aspherical $3$-manifold (a doubling argument shows these have nonpositive Euler characteristic). This answers another part of Question 1.2.
Our main tool to prove Theorems 1.3 and 1.4 is the PL Morse theory of Bestvina–Brady  [BB97]. This was inspired by Bestvina’s proof that braid groups of the tripod graph are free [Bes08, Section 4]. Discrete notions of Morse theory have been used to study graph braid groups before, however most of the earlier papers use Forman’s version [For98].
@@
We thank Ben Knudsen for comments on an earlier version of this paper.
@@ 4.3. Free product decomposition
In this section we show that if $\mathcal{P}_{m,n_{1},n_{2}}$ is the pulsar graph from Definition 4.1, then the group $B_{3}(\mathcal{P}_{m,n_{1},n_{2}})$ decomposes as a free product of the form $B_{3}(\Theta_{m})*\mathbb{F}$, where $\mathbb{F}$ is a free group. This proves Thoerem 1.4 of the introduction.